OpenStudy (tootzrll):

someone please help me on this... Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 2 h, and Car B traveled the distance in 1.5 h. Car B traveled 15 mph faster than Car A. How fast did Car B travel?

OpenStudy (irishboy123):

\(x_A = v_A \times t_A\) \(x_B = v_B \times t_B\) \(x_A = x_B \implies v_A \times t_A = v_B \times t_B\) \(t_A = 2, \; t_B = 1.5\) \(v_B = v_A + 15\) that kind of stuff :p

OpenStudy (tootzrll):

bruuuuhh im never gonna pass this... D,X

OpenStudy (irishboy123):

you know that \(distance = speed \times time\)? right?

OpenStudy (tootzrll):

yes...

OpenStudy (irishboy123):

ok so they travel the same distance, right? at different speeds and using different times.

OpenStudy (tootzrll):

yes

OpenStudy (irishboy123):

so, ..., distance travelled by car A is \(v \times 2 \) where v is the speed of the car A

OpenStudy (tootzrll):

yes

OpenStudy (irishboy123):

and that same distance, as travelled by car B, is \( (v+15)×1.5\) that might be the confusing part...

OpenStudy (tootzrll):

wait, so whats the whole equation so far? i might have it by now...

OpenStudy (irishboy123):

they travel the same distance so you can say that \[v \times 2 = (v+15)×1.5\] or \(2v = 1.5(v+15)\)

OpenStudy (tootzrll):

thanks a lot!! x)))

OpenStudy (irishboy123):

v is the speed of car A and you need the speed of car B mentioned just in case

OpenStudy (irishboy123):

\(v_B = v + 15\)